### v1

# Cross-Modal Safety Mechanism Transfer in Large Vision-Language Models

Abstract
Vision-language alignment in Large Vision-Language Models (LVLMs) successfully enables LLMs to understand visual input. However, we find that existing vision-language alignment methods fail to transfer the existing safety mechanism for text in LLMs to vision, which leads to vulnerabilities in toxic image. To explore the cause of this problem, we give the insightful explanation of where and how the safety mechanism of LVLMs operates and conduct comparative analysis between text and vision. We find that the hidden states at the specific transformer layers play a crucial role in the successful activation of safety mechanism, while the vision-language alignment at hidden states level in current methods is insufficient. This results in a semantic shift for input images compared to text in hidden states, therefore misleads the safety mechanism. To address this, we propose a novel Text-Guided vision-language Alignment method (TGA) for LVLMs. TGA retrieves the texts related to input vision and uses them to guide the projection of vision into the hidden states space in LLMs. Experiments show that TGA not only successfully transfers the safety mechanism for text in basic LLMs to vision in vision-language alignment for LVLMs without any safety fine-tuning on the visual modality but also maintains the general performance on various vision tasks (Safe and Good).
1 Introduction
Vision-language alignment methods for Large Vision-Language Models (LVLMs) use a basic LLM, a lightweight vision encoder and projector to efficiently enable the LLM to understand visual input for various vision tasks with relatively low training costs (Liu et al., 2024c; Dai et al., 2023; Zhu et al., 2023). Recent studies indicate the safety of LVLMs deserves attention (Liu et al., 2024a; Wang et al., 2023; Gong et al., 2023). Given that vision and language are aligned into a common space in LVLMs, the safety mechanism should be shared by both of them. However, this is not the case. We find that compared to toxic text input, LVLMs are more vulnerable to toxic vision input. Existing studies improve the safety of LVLMs by utilizing toxic vision data for safety fine-tuning (Zong et al., 2024; Wang et al., 2024). However, they do not provide an essential explanation for the question: “Why can’t the safety mechanism for text be shared by vision after vision-language alignment?”.
Figure 1: (a): Hidden states at the specific local transformer layers in LVLMs play a crucial role in the successful activation of safety mechanism. (b): Current vision-language alignment methods cannot effectively align vision with its semantics in text modality at hidden states level. (c): Insufficient alignment at hidden states level shifts the semantics of image and misleads the layers for safety.
Since the mainstream vision-language alignment methods for LVLMs (e.g., LLaVA (Liu et al., 2024c)) lack additional safety fine-tuning and the training data in this process contain few toxic samples (Wang et al., 2023), the safety mechanism of LVLMs is mostly inherited from the safety mechanism that have established for text in their basic LLMs. So an intuitive view to explain the dilemma of LVLMs’ safety on vision is that the vision-language alignment cannot effectively transfer the safety mechanism for text in LLMs to vision. Based on this, this paper proposes a novel perspective called Cross-Modal Safety Mechanism Transfer to rethink, explain and address the exacerbated vulnerability of LVLMs to toxic vision input compared to toxic text input. Cross-modal safety mechanism transfer means transferring the existing safety mechanism for text in LLMs to vision in vision-language alignment training without any additional safety fine-tuning on vision. Our experiments in § 3 reveal that current vision-language alignment methods fail to achieve an effective cross-modal safety mechanism transfer. LVLMs exhibit significantly different safety capabilities when handling toxicity with the same semantics but different modalities (as shown in Figure 1).
To analyze the cause of this issue, which also provides an essential explanation for the vulnerability of LVLMs to toxic vision input, we first offer an insightful understanding of where and how the safety mechanism of LVLMs operates in § 4.1. Specifically, we propose a novel method to locate the activation of safety mechanism in LVLMs and analysis the attention pattern on toxic tokens. We find that the hidden states at the specific local transformer layers play a crucial role in the successful activation of safety mechanism (Figure 1a). Then our comparative analysis between text and vision modalities in § 4.2 reveals that current vision-language alignment methods in training LVLMs cannot effectively align the vision with the language at hidden states level (Figure 1b). This misalignment causes that the hidden states of the vision shifts from the hidden states of the corresponding text with the same semantics. When the hidden states of toxic image are input to the specific transformer layers responsible for safety mechanism activation, this shift prevents these layers from correctly capturing the semantics of image as they do for text, thereby hindering their ability to accurately assess the toxicity in images (Figure 1c). This is the key reason why the safety mechanism for text in basic LLMs cannot be transferred to vision in vision-language alignment training for LVLMs.
To address the above problem, we propose a novel vision-language alignment training method called TGA, which uses retrieved texts related to the input vision to guide the projection of vision into the hidden states space in LLMs, thereby achieving alignment at hidden states level. Extensive experiments show that TGA successfully transfers the safety mechanism for text in basic LLMs to vision during vision-language alignment training for LVLMs without any safety fine-tuning on the visual modality. Besides, our TGA maintains general performance across various vision tasks compared with existing state-of-the-art (SoTA) LVLMs, such as InstructBLip, LLaVA-1.5, Qwen-VL-Chat in VQA, SEED-Bench, MM-Vet, etc. This indicates our TGA is a safe and effective vision-language alignment method for training LVLMs. The contributions of this paper are as follows:
  • We propose a novel perspective called Cross-Modal Safety Mechanism Transfer, which aims to transfer the safety mechanism for text in LLMs to vision in vision-language alignment without any safety fine-tuning on the vision data. This rethinks, explains, and addresses the exacerbated vulnerability of LVLMs to toxic vision compared to toxic text.
  • We explain where and how safety mechanism in LVLMs operates, and reveal that current vision-language alignment methods fail to achieve effective cross-modal safety mechanism transfer because of insufficient alignment at hidden states level.
  • We propose a novel vision-alignment method called TGA that can not only transfer the safety mechanism against toxic text in basic LLMs to vision but also maintain the general performance on various vision tasks compared with existing SoTA LVLMs.
2 Related Work
Vision-Language Alignment in LVLMs. Vision-language alignment in LVLMs equips basic LLMs with the ability to understand and process visual input by pre-training and instruction-tuning on large-scale text-image pairs such as works in LLaVA (Liu et al., 2024c), InstructBLip (Dai et al., 2023), Qwen-VL (Bai et al., 2023b), MiniGPT-4 (Zhu et al., 2023), Flamingo (Awadalla et al., 2023), PaLM-E (Driess et al., 2023), etc. However, it remains unknown whether visual-language alignment can extend all the capabilities of LLM on text to vision. This paper unveils that the safety mechanism of LLMs on text cannot be transferred to vision in existing typical vision-language alignment methods. We give the essential explanation for the cause of this issue and the novel solution method.
Safety of LVLMs. Recent studies have shown that LVLMs are still prone to generating toxic content when facing the toxic input (especially the toxic image), even when trained on carefully curated datasets containing few toxic samples (Wang et al., 2023; Liu et al., 2024a; Gong et al., 2023). Some studies try to address this problem by safety instruction-tuning on supervised toxic vision data (Wang et al., 2024; Zong et al., 2024). The limitations of previous studies are: (1) lacking the essential explanation for the phenomenon that safety mechanism cannot be shared by both vision and language that has been aligned. (2) collecting the multi-modal data for safety instruction-tuning is much more challenging and requires more significant human effort than only text, especially for the modalities such as audio, speech and video (Chakraborty et al., 2024). This paper gives a novel perspective that views the safety issue in LVLMs as a problem of transferring the safety mechanism for text in base LLMs to vision to rethink, explain and address the vulnerability of LVLMs to toxic vision.
3 Safety Mechanism Cannot be Cross-Modal Transferred
In this part, we provide LVLM and its basic LLM with toxic inputs having the same semantics but in different modalities (text and vision) to evaluate the safety capabilities on different modalities, which is used as an assessment of cross-modal safety mechanism transfer. We find that safety mechanism for text is not effectively transferred to vision in visual-language alignment.
Data Construction. We collect real toxic images from open-source datasets. For each image, we use LLaVA-NEXT (Liu et al., 2024b) to generate caption for it to get the toxic text-image pair. Text and image in this pair have the same semantics but are in different modalities. The specific datasets include HOD (Ha et al., 2023) that contains $10,631$ toxic images about alcohol, cigarette, gun, insulting gesture and knife, and ToViLaG (Wang et al., 2023) that contains $9,900$ toxic images about bloody and porn. After the caption generation, we get 20,31 toxic text-image pairs for experiments.
Metrics. We follow the regular safety testing method (Wang et al., 2023), which provides a toxic input and instructs the model to describe the toxic content of the input. We use Defence Success Rates (DSR), which indicates whether the model refuses to produce toxic responses when presented with toxic input, as the metric. Following (Chakraborty et al., 2024), we use LLaMA-2-7B to determine whether the responses generated by the model are toxic, thereby assessing the success of the defense.
Settings. The specific open-source LVLMs and LLMs used in experiments are: LLaVA-1.6-Mistral-7B (Liu et al., 2024b) with its basic LLM Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), InstructBlip (Dai et al., 2023) with its basic LLM Vicuna-7B-v1.5 (Zheng et al., 2024), Qwen-VL-Chat (Bai et al., 2023b) with its basic LLM Qwen-7B-Chat (Bai et al., 2023a). Some closed-source models, such as GPT-4-v, are not considered because we cannot acquire their specific checkpoints, training methods and data to provide further analysis. Given a LVLM ($\mathcal{M}$) and its basic LLM ($\mathcal{L}$), the specific settings are (1) input toxic image to $\mathcal{M}$, (2) input toxic text to $\mathcal{M}$ and (3) input toxic text to $\mathcal{L}$.
Findings. The experimental results are shown in Table 1. Different LVLMs across different toxic scenes show the following common conclusions: (1) DSR of LVLM and its basic LLM is close on toxic text, it indicates that safety mechanism of basic LLMs on text is successfully preserved in visual-language alignment training of LVLMs. (2) For the toxic information with the same semantics in different modalities (text and vision), the safety capabilities of LVLMs vary significantly, and LVLMs can hardly defense toxicity in the visual modality. It indicates that safety mechanism for text is not effectively transferred to vision in visual-language alignment training of LVLMs.
Table 1: Defence Success Rates of LVLMs and their basic LLMs for toxic input with the same semantic but in difference modalities (text and vision). “†” indicates the significant difference with p-value $\leq 0.05$ in T-test compared with the toxic text input.
Scenes LLaVA-1.6-Mistral-7B InstructBlip Qwen-VL-Chat
Basic LLM LVLM Basic LLM LVLM Basic LLM LVLM
Toxic Text Toxic Text Toxic Image Toxic Text Toxic Text Toxic Image Toxic Text Toxic Text Toxic Image
Porn 27.45 26.78 1.05 17.98 15.46 1.28 72.96 70.64 4.23
Bloody 12.31 11.72 0.56 8.46 7.11 0.12 35.60 33.86 1.46
Insulting Gesture 43.18 42.97 0.78 30.75 30.22 0.57 78.74 76.26 5.15
Alcohol 32.52 32.03 0.25 21.60 21.05 0.00 85.05 83.25 5.48
Cigarette 22.57 22.84 0.17 15.81 15.76 0.00 83.76 82.80 4.41
Gun 52.46 51.94 1.95 39.45 39.27 0.75 86.42 86.90 5.72
Knife 45.91 43.06 1.22 30.77 30.60 0.23 89.01 88.73 5.40
4 Cause of Failure in Cross-Modal Transferring Safety
This section analyzes the cause of the failure in transferring the safety mechanism from text to vision. Firstly, in § 4.1, we find that hidden states at the specific transformer layers play a crucial role in the successful activation of safety mechanism. Then, our comparative analysis between text and image in § 4.2 reveals that alignment between vision and language at hidden states level is insufficient. It makes the transformer layers responsible for safety mechanism activation cannot correctly capture the semantics of image as they perform on text, so they cannot correctly assess the toxicity in image and the safety mechanism on vision collapses.
4.1 Safety Mechanism is Activated at Specific Layers by Hidden States
This part gives an insightful explanation of where and how the safety mechanism activated in LVLMs on textual, which is fundamental to understand the safety mechanism collapse on vision. Previous studies (Tenney, 2019; Dai et al., 2021; Meng et al., 2022) have shown that transformers in different layers of language model have different functions such as lexical, semantic, knowledge, etc., this paper proposes a novel method to identify the transformer layers responsible for activating safety mechanism in LVLMs and analyzes the attention patterns on toxic tokens within these layers to give the explanation of where and how the safety mechanism is activated.
Where: Locating the Activation of Safety Mechanism. When facing a toxic input, if the safety mechanism is activated, the language model will refuse to follow the instruction and apologize to the user, such as "Sorry but I cannot …", otherwise, the language model will follow the instruction and generate the corresponding response (Bianchi et al., 2024). Therefore, a reply with an apology is an important signal for the activation of the safety mechanism. This paper proposes a novel method to locate where the safety mechanism is activated by detecting the word distribution mutation on sorry semantics among layers. Specifically, given a toxic text $t$ and instruction $s$ to LVLMs, the next token prediction for $x$ can be formalized as:
$\textrm{P}(x|t,s)=\textrm{softmax}(\textrm{{W}}\textrm{{H}}^{\prime}),x\in\mathcal{X},$ (1)
in which $\textrm{{W}}\in\mathbb{R}^{h\times v}$ is the vocabulary head that maps the output hidden states $\textrm{{H}}^{\prime}$ to the word distribution in vocabulary $\mathcal{X}$ with size $v$. Since previous studies (Chuang et al., 2024; Xu et al., 2024) prove that due to the residual connections, the combination of any intermediate layer in a language model with the vocabulary head W can represent its distribution of semantics, we apply W to each intermediate layer to obtain its distribution of semantics as:
$\textrm{P}_{j}(x|t,s)=\textrm{softmax}(\textrm{{W}}\textrm{{H}}_{j}),x\in\mathcal{X},$ (2)
in which $j$ is the $j\mbox{-}th$ transformer layer of LVLMs, $\textrm{{H}}_{j}$ is the hidden states of the last token in $j\mbox{-}th$ layer. We calculate the word distribution change in $j\mbox{-}th$ layer by contrasting the $(j-1)\mbox{-}th$ layer as:
$\textrm{D}_{j}(x|t,s)=\textrm{log}\frac{\textrm{P}_{j}(x|t,s)}{\textrm{P}_{j-1}(x|t,s)},j>1.$ (3)
Algorithm 1 Find the layer where safety mechanism is activated.
1:for $j=2$ to $N$ do $\triangleright$ Traverse word distribution change from 1 to N transformer layers.
2:     if $\operatorname*{arg\,max}\textrm{D}_{j}(x|t,s)$in $\mathcal{K}$
$\triangleright$ Find the layer where sorry tokens start to rank Top-1 in word distribution change.
3:         return $j$
4:     end if
5:end for
6:return $-1$triangleright If no such layer is found, return $-1$
We collect the tokens with sorry semantics such as “sorry”, “apologize" from vocabulary $\mathcal{X}$ and denote them as set $\mathcal{K}$. We add prompt that instructs LVLMs to response as "Sorry but I cannot…" if the input contains toxicity. Then, we propose Algorithm 1 to locate the layer where safety mechanism is activated by finding the layer where sorry tokens start to rank Top-1 in word distribution change.
Sorry tokens start to show the most significant growth among the whole vocabulary means that the transformer layer realizes that the current input is toxic, so it tries to update the word distribution to refuse following the instruction, which is actually the safety mechanism activation. So our method makes sense. Safety mechanism perturbation experiments in Figure 3 also show that our method can accurately determine the activation location of the safety mechanism of LVLMs.
(a) LLaVA-1.6-Mistral-7B
(b) InstructBlip
(c) Qwen-VL-Chat
Figure 2: Location of safety mechanism activation and attention map on toxic tokens varies with layer in LVLMs. The blue line is the proportion $R$ of attention from token $x$ to the toxic token set $\mathcal{C}$ over the entire attention map varies with layer. The pink region are the layers where safety mechanism is activated. The left and right boundaries of the region are the minimum and maximum activation layer on the entire sample set respectively.
(a) LLaVA-1.6-Mistral-7B
(b) InstructBlip
(c) Qwen-VL-Chat
Figure 3: Defense Success Rate (DSR) for the toxic text input in the condition that information flow from toxic tokens are masked in every $5$ layers. The yellow bars are the original DSR without any truncating. The pink bars are truncating at layers where safety mechanism is activated determined by our method. The blue bars are truncating at other layers.
How: Analysing the Attention Patterns on Toxic Tokens. For each toxic text in Data Construction of § 3, we use gpt-4o-2024-05-13 API to extract the specific toxic words in the text and mark them as the toxic tokens (denoted as the set $\mathcal{C}$) for LVLM’s input. In the word distribution prediction of $\textrm{P}_{j}(x|t,s)$ in the $j\mbox{-}th$ layer with toxic text $t$ and instruction $s$ as the input, we obtain the attention map $\mathcal{A}_{j}$ of the last token to tokens in $t$, and calculate the proportion $R$ of attention score from the last token to the toxic token set $\mathcal{C}$ over the entire attention map $\mathcal{A}_{j}$ as:
$\mathcal{A}_{j}=\text{softmax}\left(\frac{Q_{j}K_{j}^{\top}}{\sqrt{d_{k}}}\right),R=\sum_{i\in\mathcal{C}}\mathcal{A}_{j}^{i},$ (4)
in which $Q_{j}\in\mathbb{R}^{1\times d_{k}}$ is the query vector of the last token, $K_{j}\in\mathbb{R}^{n\times d_{k}}$ is the key vector for the input text $t$, where $n$ denotes the number of tokens in the input text $t$.
Practical Explanation of Safety Mechanism Activation. Based on the methods in above two parts about identifying where and how the safety mechanism is activated in LVLMs, we select samples that LVLMs successfully execute safety mechanism on the toxic text input from the dataset of § 3 and analyze the activation location and attention patterns to provide a practical explanation for the activation of safety mechanism. The average statistical results of the above two points on the entire sample set are shown in Figure 2. They indicate that the activation of the safety mechanism coincides with peaks in attention to the hidden states of toxic tokens. Therefore, a conclusion can be obtained that the safety mechanism is activated by the information from the hidden states of toxic tokens at the specific transformer layers. Following Safety Mechanism Perturbation experiments support this.
Safety Mechanism Perturbation. To support above conclusion, we conduct corresponding safety mechanism perturbation experiments in which the information from toxic tokens is masked at different layers, and test the changes in the safety capabilities of LVLMs under these conditions. We achieve this by adding mask to toxic tokens in attention map in specific layers. The datasets are also the same as Table 1 and we use the average DSR on seven scenes as the metric. Experimental results in Figure 3 show that truncating at the layers where safety mechanism is activated determined by our method causes the most significant disruption to the safety mechanism of LVLMs compared to other layers. This further demonstrates the effectiveness of our method and the validity of our conclusion.
4.2 Insufficient Alignment at Hidden States Misleads Safety Mechanism
This section gives our analysis about the cause of the failure in transferring the safety mechanism from text to vision in LVLMs. As our findings in § 4.1, the hidden states of input tokens at specific transformer layers play a crucial role in the successful activation of safety mechanism. We conduct the comparative analysis between hidden states of input text and image with the same semantics, which helps us find the essential explanation for the failure of cross-modal safety mechanism transfer.
Comparative Analysis Between Text and Image. Figure 4 is the cosine similarity between the mean pooled vectors of hidden states of texts and images with the same semantics. At the layers where safety mechanism is activated, compared with the Clip Score that indicates the semantic similarity between text and image, the cosine similarity between hidden states of text and image in LVLMs is significantly lower, suggesting that the alignment between text and image at the hidden states level is insufficient. The transformer layers for safety mechanism activation cannot correctly capture the semantics of image as they perform on text, so they cannot correctly assess the toxicity in image.
(a) LLaVA-1.6-Mistral-7B
(b) InstructBlip
(c) Qwen-VL-Chat
Figure 4: Cosine similarity between hidden sates of text and image input with the same semantics varies with layer in LVLMs. The pink region are the layers where safety mechanism is activated on text. The left and right boundaries of the region are the minimum and maximum activation layer on the entire sample set respectively. The purple dashed line is the cosine similarity between the output representations of text-image pairs obtained by CLIP ViT-H/14.
(a) LLaVA-1.6-Mistral-7B
(b) InstructBlip
(c) Qwen-VL-Chat
Figure 5: Defense Success Rate (DSR) for the toxic image input in the condition that hidden states in every $5$ layers are added by mean pooled hidden states of the text. The yellow bars are the original DSR without any operation. The pink bars are adding in layers where safety mechanism is activated determined by our method. The blue bars are adding in other layers.
Safety Mechanism Reconstruction. The above analysis suggests that the insufficient alignment of hidden states between vision and language in the layers where safety mechanism activation is a potential reason for the safety mechanism collapse on vision. We perform safety mechanism reconstruction experiments to demonstrate this. Specifically, for the input image $I$, its hidden states sequence for image tokens at the $j\mbox{-}th$ transformer layer in LVLMs is denoted as ${\mathbb{I}}_{j}=\{\textrm{{I}}^{1}_{j},\textrm{{I}}^{2}_{j},...,\textrm{{I}}^{n}_{j}\}$, in which $\textrm{{I}}^{k}_{j}$ is the hidden state for the $k\mbox{-}th$ image token, $n$ is the number of image tokens. Similarly, for the text $T$ having the same semantics with image $I$, its hidden states sequence for text tokens at the $j\mbox{-}th$ layer is ${\mathbb{T}}_{j}=\{\textrm{{T}}^{1}_{j},\textrm{{T}}^{2}_{j},...,\textrm{{T}}^{m}_{j}\}$. We adjust ${\mathbb{I}}_{j}$ as:
${\mathbb{I}}_{j}=\{\textrm{{I}}^{k}_{j}+\textrm{mean}(\displaystyle{\mathbb{T}}_{j})\mid\textrm{{I}}^{k}_{j}\in\displaystyle{\mathbb{I}}_{j}\},\textrm{mean}(\cdot)\textrm{ is mean pooling.}$. (5)
By adding the mean pooled hidden states of text $T_{j}$ to $I_{j}$ as the input for the $(j+1)\mbox{-}th$ layer, the hidden states of image $I$ are forcibly aligned with the corresponding text $T$. This operation is performed on different transformer layers and the corresponding safety capabilities on toxic image input are shown in Figure 5. Directly adding hidden states of text to hidden states of image can significantly improve the safety capabilities of LVLMs on toxic image, especially operating on safety mechanism activation layers determined by our method, this further indicates that the alignment between vision and language at hidden states level plays a critical role in the successful transfer of the safety mechanism from text to vision in LVLMs.
Explanation of Safety Mechanism Collapse in Transfer from Text to Vision. Based on § 4.1 and § 4.2, the following conclusions can be obtained to explain why the safety mechanism existing in basic LLMs for text cannot be cross-modal transferred to vision in vision-language alignment:
  • (1) Safety Mechanism is activated at the specific transformer layers.
  • (2) Hidden states at the specific transformer layers play a crucial role in the successful activation of safety mechanism.
  • (3) Existing vision-language alignment methods for LVLMs fail to align the hidden states of vision with its corresponding hidden states of text, especially in the transformer layers that responsible for activating safety mechanism.
  • (4) This misalignment makes the transformer layers responsible for safety mechanism activation cannot correctly capture the semantics of image as they perform on text, so they cannot correctly assess the toxicity in image and the safety mechanism on vision collapses.
5 Text Guided Alignment at Hidden States Level
Our analysis in § 4 indicates that insufficient alignment of hidden states between vision and language in the layers where safety mechanism activation is a critical reason for the safety mechanism collapse in transfer from text to vision. To address this, this section proposes a novel text-guided vision-language alignment method that can effectively align vision and language at hidden states level.
5.1 Motivation and Overview
As shown in Figure 6 (a), previous vision alignment methods in LVLMs construct image-instruction-output triples as training data, and use cross-entropy in language modeling as the loss fuction to optimize the output generated by basic LLMs to make LLM understand the input vision, thereby achieving visual-language alignment. We name these methods input-to-output alignment, i.e. align the text output to the visual input, which is actually asymmetrical. The natural flaw of these method is that they black-box basic LLMs and only focus on output, while neglecting whether the internal representation of the visual input in LLMs, i.e., the hidden states, aligns with the hidden states in text modality. To address this problem, we propose the Text-Guided input-to-input Alignment (TGA) at hidden states level, as shown in Figure 6 (b). For the input image, TGA retrieves the semantically relevant text as the template to guide the alignment of vision to language at hidden states level.
Figure 6: Comparison between previous methods and our TGA.
5.2 Detailed Method
5.2.1 Data Construction
The original training data is collected from LLaVA (Liu et al., 2024c) that consists of $558$K images for pre-training and $665$K images for instruction-tuning. Each data sample is in multi-turn conversation template like this: ($\textbf{X}_{\textrm{image}}$, $\textbf{X}_{\textrm{inst}}^{1}$- $\textbf{X}_{\textrm{r}}^{1}$, $\textbf{X}_{\textrm{inst}}^{2}$-$\textbf{X}_{\textrm{r}}^{2}$, …, $\textbf{X}_{\textrm{inst}}^{q}$-$\textbf{X}_{\textrm{r}}^{q}$). $\textbf{X}_{\textrm{image}}$ is the input image, $\textbf{X}_{\textrm{inst}}$-$\textbf{X}_{\textrm{r}}$ pair is one conversation turn consisting of an instruction $\textbf{X}_{\textrm{inst}}$ and its response $\textbf{X}_{\textrm{r}}$. For each image $\textbf{X}_{\textrm{image}}$ in the training set, our method uses a image-to-text retrieval model BEIT-3 (Wang et al., 2022) to retrieve a relevant text $\textbf{X}_{\textrm{retrieval}}$ from a large-scale corpus that consists of $1,153$K textual captions for images in LAION/CC/SBU (Schuhmann et al., 2021; Changpinyo et al., 2021) dataset with little toxicity, and uses LLaVA-1.5-13B to generate a textual caption $\textbf{X}_{\textrm{caption}}$ for $\textbf{X}_{\textrm{image}}$, which can be seen as the description of image semantics in text modality.
5.2.2 Training
At each training step, TGA firstly inputs $\textbf{X}_{\textrm{caption}}$ to LVLMs to get the hidden states of $\textbf{X}_{\textrm{caption}}$ at each transformer layer under the condition of disabling gradient calculation:
${\mathbb{C}}_{j}=\{\textrm{{C}}^{1}_{j},\textrm{{C}}^{2}_{j},...,\textrm{{C}}^{m}_{j}\},j=1,2,3,...,N,$ (6)
[…]
$\mathcal{L}_{\textrm{guide}}=\sum_{j=1}^{N}-\textrm{cos}(\displaystyle\overline{{\mathbb{I}}_{j}},\displaystyle\overline{{\mathbb{C}}_{j}})+\log\left[1+\underbrace{\exp\left[-(\textrm{cos}(\displaystyle\overline{{\mathbb{I}}_{j}},\displaystyle\overline{{\mathbb{C}}_{j}})-\textrm{cos}(\displaystyle\overline{{\mathbb{R}}_{j}},\displaystyle\overline{{\mathbb{C}}_{j}}))\right]}_{\textbf{Pair-wise}}\right],$ (7)
in which $\overline{{\mathbb{I}}_{j}}$, $\overline{{\mathbb{C}}_{j}}$ and $\overline{{\mathbb{R}}_{j}}$ are mean pooled vectors of hidden states of $\textbf{X}_{\textrm{image}}$, $\textbf{X}_{\textrm{caption}}$ and $\textbf{X}_{\textrm{retrieval}}$ in the $j$-$th$ transformer layer respectively. The intuition of this pair-wise loss is to ensure that $\textbf{X}_{\textrm{image}}$ does not directly replicate the semantics of $\textbf{X}_{\textrm{retrieval}}$, but rather uses $\textbf{X}_{\textrm{retrieval}}$ as a template for partially similar semantics in the text modality to prompt the alignment of ${\mathbb{I}}_{j}$ with its text modality hidden states ${\mathbb{C}}_{j}$. The successful alignment should achieve that ${\mathbb{I}}_{j}$ is closer to ${\mathbb{C}}_{j}$ than ${\mathbb{R}}_{j}$, because in this condition, ${\mathbb{I}}_{j}$, ${\mathbb{C}}_{j}$ and ${\mathbb{R}}_{j}$ are aligned into a common space and ${\mathbb{C}}_{j}$ and ${\mathbb{I}}_{j}$ have consistent semantics. $\textrm{cos}(\displaystyle\overline{{\mathbb{R}}_{j}},\displaystyle\overline{{\mathbb{C}}_{j}})$ is used as the lower bound supervision for alignment between vision $\overline{{\mathbb{I}}_{j}}$ and language $\overline{{\mathbb{C}}_{j}}$. The total loss function $\mathcal{L}$ is the combination of $\mathcal{L}_{\textrm{guide}}$ and cross-entropy loss for language modeling:
$\mathcal{L}$ $=\mathcal{L}_{\textrm{guide}}-\frac{1}{N}\sum_{i=1}^{N}\log\textrm{P}(\textbf{X}_{\textrm{a},i}|\textbf{X}_{\textrm{retrieval}},\textbf{X}_{\textrm{image}},\textbf{X}_{\textrm{inst}},\textbf{X}_{\textrm{a},<i}),\textbf{X}_{\textrm{a}}\textrm{ is the answer for }\textbf{X}_{\textrm{inst}}.$ (8)
5.3 Experiments
5.3.1 Experimental Details
Experimental Setting. Our method aims to transfer the safety mechanism of LLMs on text to vision during the vision-language alignment in LVLMs, so the main setting in this experiment is assessing the safety capabilities of LVLMs on toxic image without any safety fine-tuning on visual modality. We follow the regular safety testing method (Wang et al., 2023) that gives a toxic image and instructs the model to describe the toxic information of the image. We use Defence Success Rates (DSR) that indicates the whether a model refuse to produces toxic responses when presented with the toxic image as the metric. We follow (Chakraborty et al., 2024) to use LLaMA-2-7B to determine whether the responses generated by the model are toxic, that is, whether the defense is successful.
Datasets. The training datasets for TGA are consistent wiht LLaVA (Liu et al., 2024c) that collects 558K images for pre-traing and 665K images for instruction-tuning. The pre-training dataset is a filtered subset of CC3M and the instruction-tuning dataset is LLaVA-1.5-mix665k. Datasets to evaluate the safety capabilities on vision are consistent with Section 3, which include $20,531$ toxic images about alcohol, cigarette, gun, insulting gesture, knife, bloody and pornography.
Baselines. Baselines in this experiments can be categorized into two types. The first type is exiting mainstream state-of-the-art vision-language alignment training method such as LLaVA (Liu et al., 2024b), InstructBlip (Dai et al., 2023) and Qwen-VL (Bai et al., 2023b). The second type is Unlearn-FigS (Chakraborty et al., 2024), a method that performs textual unlearning on VLMs to enhance the cross-modal safety capabilities. Since our paper introduces a pioneering concept of transferring safety mechanism for text to vision without safety fine-tuning on vision, while existing methods for the safety of LVLMs, such as (Zong et al., 2024), rely on supervised toxic vision data for safety fine-tuning. Therefore, these methods are unfair in our setting and are not considered.
Implementation Details. We use Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) as basic LLM, clip-vit-large-patch14-336 (Radford et al., 2021) as the vision tower and two-layer MLP as projector. Our model is training on 64 × V100 GPUs with Deepspeed Zero-Stage 3 as acceleration framework in float32. Most hyperparameters follow LLaVA (Liu et al., 2024c). In pre-training, we freeze basic LLMs and only train projector for 1 epoch with a learning rate of 2e-3. In instruction-tuning, we make all parameters trainable with a learning rate of 2e-6 for 1 epoch. As for the image-to-text retrieval, our retrieval database consisting of $1,158$K texts, in which $1,153$K texts are captions for images in LAION/CC/SBU dataset with little toxicity, filtered with a more balanced concept coverage distribution, and 5K texts are toxic texts across various harmful domains generated by Vicuna-13B. We use BEIT-3 (Wang et al., 2022) to index these texts and retrieve the top-1 text for each image.
5.3.2 Experimental Results
Table 2: Defence Success Rates on toxic image across seven scenes in different vision-language alignment training method in LVLMs. Safety of LLM is referred to Table 1.
Method Basic LLM Safety of LLM Defence Success Rates on Toxic Scenes
Porn Bloody Insulting Alcohol Cigarette Gun Knife
BLIP-2 Vicuna-13B Weak 1.17 0.00 0.00 0.00 0.00 0.98 0.06
InstructBlip Vicuna-7B Weak 1.28 0.12 0.57 0.00 0.00 0.75 0.23
LLaVA-1.5 Vicuna-7B Weak 1.20 0.37 0.57 0.19 0.76 1.22 0.35
LLaVA-1.6 Mistral-7B Medium 1.05 0.56 0.78 0.25 0.17 1.95 1.22
Qwen-VL-chat Qwen-7B Srong 4.23 1.46 5.15 5.48 4.41 5.72 5.40
Unlearn-FigS Mistral-7B Medium 8.76 4.27 16.98 14.31 10.10 21.42 18.55
TGA (Ours) Mistral-7B Medium 20.65 9.48 22.73 17.92 17.29 30.83 29.42
Main Results. The defence success rates on toxic image across seven scenes in different vision-language alignment training method in LVLMs are shown in Table 2. Our TGA significantly improve the safety capabilities of LVLMs on toxic image than mainstream vision-language alignment method, without any additional safety fine-tuning on vision. It is because that our TGA successfully transfers the safety mechanism present in basic LLMs for text to vision by improving the vision-language alignment at hidden states level. Besides, our TGA outperforms Unlearn-FigS, a method that works by transferring text unlearning from LLMs to LVLMs to avoid harmful context toward toxic regions, which shows that directly transferring safety mechanism in our TGA is a more effective method.
General Capabilities of LVLMs on Vision Tasks. Table 3 shows that our TGA shows comparable performance on various vision tasks. Combined with Table 2, our TGA is a safe and good vision-language alignment method for training LVLMs, it not only successfully transfers the safety mechanism from text to vision but also maintains the general performance on various vision tasks. The training data of LLaVa-1.6 has not been open but is more effective than the data in LLaVA-1.5 used by our method, therefore, to keep the fair comparison, we reproduce LLaVA-1.5 on Mistral to replace it. The specific details about benchmarks and metrics in Table 3 can be found in Appendix A.1.
Table 3: Comparison with SoTA methods on benchmarks evaluating LVLMs on various vision tasks.
Method SciQA POPE SEED-Bench MM-Vet
img-acc rand pop adv all img vid rec ocr know gen spat math all
BLIP-2-13B 61.0 89.6 85.5 80.9 46.4 49.7 36.7 - - - - - - 22.4
InstructBlip-7B 60.5 - - - 53.4 58.8 38.1 - - - - - - 26.2
InstructBLIP-13B 63.1 87.7 77 72 - - - - - - - - - 25.6
Qwen-VL-chat-7B 68.2 - - - 58.2 65.4 37.8 36.8 21.9 16.3 19.7 23.5 6.3 30.2
LLaVA-1.5-7B 66.8 87.3 86.1 84.2 58.6 66.1 37.3 37.0 21.0 17.6 20.4 24.9 7.7 31.1
LLaVA-1.5-7B (Mistral) 67.9 87.3 86.2 84.0 58.8 66.5 37.4 37.2 22.5 20.1 22.3 28.6 7.7 32.9
TGA-7B (Ours) 71.7 87.6 86.2 83.7 58.7 66.3 37.4 37.7 28.9 23.0 25.6 37.6 7.7 34.6
(a) Safety mechanism transfer rates from language to vision.
(b) Cosine similarity between hidden states of texts and images with the same semantics.
(c) Ablation study about the effect of retrieved text.
Figure 7: Analysis and ablation study.
Safety Mechanism Transfer. Figure 7(a) shows that compared with vision-language alignment methods in baselines, our TGA effectively transfer the safety mechanism from language to vision. In this experiment, we evaluate the DSR of basic LLMs on toxic text and the DSR of aligned LVLMs on toxic image. The ratio between the two serves as the safety transfer rate from language to vision.
Alignment at Hidden States Level. Figure 7(b) shows that comapred with the baseline, our TGA effectively aligns vision and language at hidden states level. Since the training data and hyperparameter in our TGA is consistent with LLaVA-1.5, so we use LLaVA-1.5 based on Mistral-7B as the baseline. Our TGA achieves the greater cosine similarity between hidden states of texts and images with the same semantics than the baseline.
Ablation Study on Retrieved Text. The main ablation study of our TGA is about the introduction of retrieved text. As shown in Figure 7(c), we explore the effect of retrieved text when used in training and inference. For baseline LLaVA-1.5, the usage of retrieved text in inference cannot significantly improve its safety capabilities. For our TGA, even if without retrieved text, our method can still outperforms LLaVA-1.5 in safety. The usage of retrieved text with loss $\mathcal{L}_{\textrm{guide}}$ in Equ. 7 can further improve the safety capabilities and general performance on vision tasks.
Case Study. We show the case study about baselines and our methods facing toxic input in different scenes. It can be found in Appendix A.2.
6 Conclusion and Discussion
This paper proposes a novel perspective called Cross-Modal Safety Mechanism Transfer to rethink, explain and address the exacerbated vulnerability of LVLMs to toxic vision compared to toxic text. Extensive analysis shows that current vision-language alignment methods fail to achieve effective cross-modal safety mechanism transfer and the reason is the insufficient alignment between vision and language at hidden states level. To address this, we propose a novel vision-language alignment method that can not only transfer the safety mechanism against toxic text in basic LLMs to vision but also maintain the general performance on various vision tasks compared with existing SoTA LVLMs.
Appendix A More Detailed Experiments
A.1 General Capabilities of LVLMs on Vision Tasks
Datasets, Benchmarks and Metrics.
(1) ScienceQA (Lu et al., 2022): A benchmark that consists of  21k multimodal multiple choice questions with a diverse set of science topics and annotations of their answers with corresponding lectures and explanations. We follow LLaVA (Liu et al., 2024c) to evaluate the zero-shot generalization of LVLMS on scientific question answering in image subset and use accuracy as the metric.
(2) POPE (Li et al., 2023b): POPE evaluates model’s degree of hallucination on three sampled subsets of COCO (Lin et al., 2014): random, common, and adversarial and we report the F1 score as the metric on all three splits.
(3) SEED-Bench (Li et al., 2023a): SEED-Bench consists of 19K multiple choice questions with accurate human annotations (x 6 larger than existing benchmarks), which spans 12 evaluation dimensions including the comprehension of both the image and video modality. We report the accuracy on image, video and all as the metric.
(4) MM-Vet (Yu et al., 2023): MM-Vet evaluate model’s capabilities in engaging in visual conversations on a diverse range of tasks, and evaluates the correctness and the helpfulness of the response with GPT-4 evaluation.
A.2 Case Study
Case study about response of different LVLMs for toxic image is shown in Figure 8 to 14.
Figure 8: Case study on toxic input image about alcohol.
Figure 9: Case study on toxic input image about bloody.
Figure 10: Case study on toxic input image about cigarette.
Figure 11: Case study on toxic input image about gun.
Figure 12: Case study on toxic input image about porn.
Figure 13: Case study on toxic input image about insulting gesture.
Figure 14: Case study on toxic input image about knife.